Good morning!
I have to find the inverse of the function: f(x)= 4x+1/3.
Thanks so much for all your help!
Denise

Answer by Jk22(389)

Let write the inverse g(x)=f^(-1)(x), then we should have :
f(g(x)) = x = 4g(x)+1/3 => g(x) = (x-1/3)/4

we can see that g(f(x)) = (f(x)-1/3)/4 = (4x+1/3-1/3)/4 = x

Note : the functional inverse is not the inverse in the real numbers, which were denoted f(x)^(-1) = 1/(4x+1/3)